# Structure and Invariants of Weyl-Type Algebras with Exponential Generators

### v1

@@ 5 Center and Central Simplicity of Weyl-Type Algebras
Since $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over $Z(A)=\mathbb{F}[y^{\pm 1}]$, it defines an element in $\operatorname{Br}(\mathbb{F}(y))$ (the Brauer group of the function field). The degree of $A$ over its center is $2^{\operatorname{rank}(\mathcal{A})}$, so its period divides $2^{\operatorname{rank}(\mathcal{A})}$ by general theory of central simple algebras.
To see nontriviality, note that the specialization at $y=1$ gives the classical Weyl algebra with exponential generators, which is known to be nontrivial in the Brauer group. ∎
@@
    The center is $Z(A)=\mathbb{F}[e^{\pm x^{p}e^{t}}]$, never equal to $\mathbb{F}$ (nontrivial).
@@
    $A$ defines a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ where $y=e^{x^{p}e^{t}}$.
@@ 6 Open Problems
The study of non-associative algebras with exponential generators, their Gelfand–Kirillov dimension, automorphism groups, and central structure leads naturally to several unresolved questions. Below we formulate two open problems that arise directly from the results presented in this paper.
Open Problem 1: GK Dimension of Non-associative Algebras with Exponential Generators
@@
Problem 6.1 (GK Dimension of Non-associative Exponential Algebras).
@@
$N=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
@@
  • How does $\operatorname{GKdim}(N)$ compare with the GK dimension of its associative envelope $A\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$?
@@
$N_{p,t}=N\!\left(\mathbb{F}[e^{\pm x^{p}e^{t}},\;e^{\mathcal{A}x},\;x^{\mathcal{A}}]\right)$
@@
This problem connects with the broader study of deformations of algebraic structures HartwigLarssonSilvestrov06 and the classification of non-associative algebras Schafer17. A solution would deepen our understanding of how non-associativity affects the moduli space of exponential Weyl-type algebras.
@@
  • Physical applications: Such algebras appear in certain deformed quantum mechanical models. A systematic study of their physical interpretations is yet to be carried out.
Progress on these open problems would not only advance the theory of non-associative algebras with exponential generators but also provide new insights into the interplay between non-associativity, growth, and deformation theory.

### v2

@@ 5 Center and Central Simplicity of Weyl-Type Algebras
Since $A_{p,t,\mathcal{A}}$ is an Azumaya algebra over $Z(A)=\mathbb{F}[y^{\pm 1}]$, it naturally defines an element in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ of the function field $\mathbb{F}(y)$. The degree of $A$ over its center is $2^{\operatorname{rank}(\mathcal{A})}$, so by the general theory of central simple algebras, its period divides $2^{\operatorname{rank}(\mathcal{A})}$.
To see nontriviality, consider the specialization at $y=0$ (or any generic value). The resulting algebra is a twisted Weyl algebra with hyperbolic sine generators, which is known to represent a nontrivial class in the Brauer group. Hence $[A_{p,t,\mathcal{A}}]$ is nontrivial. ∎
@@
    The center is $Z(A)=\mathbb{F}[\sinh(\pm x^{p}\sinh(t))]$, never equal to $\mathbb{F}$ (nontrivial).
@@
    $A$ defines a nontrivial class in the Brauer group $\operatorname{Br}(\mathbb{F}(y))$ where $y=\sinh(x^{p}\sinh(t))$.
@@ 6 Open Problems
The study of non-associative algebras with hyperbolic sine generators, their Gelfand–Kirillov dimension, automorphism groups, and central structure naturally leads to several unresolved questions. Below we formulate two open problems that arise directly from the results presented in this paper.
Open Problem 1: GK Dimension of Non-associative Algebras with Hyperbolic Sine Generators
@@
Problem 6.1 (GK Dimension of Non-associative Hyperbolic Sine Algebras).
@@
$N=N\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$
@@
  • How does $\operatorname{GKdim}(N)$ compare with the GK dimension of its associative envelope $A\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$?
@@
$N_{p,t}=N\!\left(\mathbb{F}[\sinh(\pm x^{p}\sinh(t)),\;\sinh(\mathcal{A}x),\;x^{\mathcal{A}}]\right)$
@@
This problem connects with the broader study of deformations of algebraic structures HartwigLarssonSilvestrov06 and the classification of non-associative algebras Schafer17. A solution would deepen our understanding of how non-associativity affects the moduli space of hyperbolic sine Weyl-type algebras.
@@
  • Physical applications: Such algebras appear in certain deformed quantum mechanical models involving hyperbolic potentials. A systematic study of their physical interpretations is yet to be carried out.
Progress on these open problems would not only advance the theory of non-associative algebras with hyperbolic sine generators but also provide new insights into the interplay between non-associativity, growth, and deformation theory.